Solving systems of linear equations
Alesha is 24 years older than Zack. In 13 years Alesha will be double Zack age. How old are Alesha and Zack?

Answer by ikleyn(52803)

.
Solving systems of linear equations
Alesha is 24 years older than Zack. In 13 years Alesha will be double Zack age. How old are Alesha and Zack?
~~~~~~~~~~~~~~~~~~~~~~~~~~~

A = Z + 24          (1)     ("Alesha is 24 years older than Zack.")
A + 13 = 2(Z+13)    (2)     ("In 13 years Alesha will be double Zack age")

You just have A expressed via Z in (1): A = Z + 24.
Substitute it into (2) replacing A. You will get

(Z + 24) + 13 = 2Z + 26.

Solve for Z.


Answer.  Z = 11,  A = 35.

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - OVERVIEW of lessons on age problems
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".